Thursday, Mar. 03: OWL Shift

From Hall A Wiki
Jump to: navigation, search

Go back to Daily Run Plan Daily Run Plans

General considerations

  • BTA: SL makes sure to submit his/her BTA successfully and be shown in HALOG
  • Need shift crews for future

Specific mission

  • Production Run
  • Beam set up without Compton chicane for tonight (will try to set beam up with Compton chicane after RF recovery and BCM calibration tomorrow afternoon)

Upcoming event(Short Term)

  • Note: MCC RF recovery (8:30-11:30AM, Thursday)

Access Wish List:(controlled access)
1/ Doug: Series of magnet measurement <: 1hr / no beam/ but beam permit
(a) He will start 8:30AM at MCC
(b) Expect to complete by 9:00AM (or whenever he finishes, he will notify us.)

After RF recovery,

  • Mar/03 BCM calibration (high beam current 30uA(60 ideal),FC, invasive Hall-D after RF recovery done (around noon)

-- Hall-D will be down for adding new diamond radiator (up to 8hrs): This is good chance for us to BCM calibration

  • Detailed instructions


  • If everything goes well, We come back to production run

Upcoming event(Long Term)

  • Target boiling test
  • Beam study (Thursday, 3/10)

->Then call MCC for controlled access
Since short break (RF recovery) and not crucial access tomorrow (and people agree that),
The controlled access will be postponed till next beam study which will be Thursday next week.

2/ Andrew Lumanog: (Call 327-5523) : <1hr
3/ Dave/Greg: check Compton laser profile : <1hr
4/ Longwu/Thir: D1 BCM for RHRS

BPM/Spot++ Check

  • After beam lost for considerable time (like beam study, pass change,...)
  • Ask MCC to run 5uA wth carbon hole target and make sure you can see the hole through Spot++